OpenStudy (anonymous):

Quadrilateral GRAY below is an isosceles trapezoid with bases GR and AY. If m∠G = 7x – 3 and m∠R = 3x + 33, what is the measure of ∠G? A. 9° B. 15° C. 60° D. 78°

OpenStudy (anonymous):



OpenStudy (anonymous):

2(3x + 33) + 2 (7x – 3) = 360

OpenStudy (anonymous):

6x + 66 + 14x + 6 = 360

OpenStudy (anonymous):

72 + 20x = 360

OpenStudy (anonymous):

20x = 360 - 72 20x = 288

OpenStudy (anonymous):

288/20 = x

OpenStudy (anonymous):

x = 14.4

OpenStudy (anonymous):

theres no answer choice thats 14.4, are you sure?